OpenStudy (anonymous):

What are the 3 requirements a function must meet in order to be able to use Rolle's Theorem? I know the first 2 are "Is the function continuous?" and "Is the function differentiable?", what's the 3rd requirement?

OpenStudy (freckles):

look like the if part of rolle's theorem that which is in the if part must be the part in which the function meets as a requirement to apply rolle's theorem

OpenStudy (freckles):

If I is continuous on a closed interval [a,b] and differentiable on the open interval (a,b) and if f(a)=f(b), then there is at least one point c in (a,b) such that f'(c)=0.

OpenStudy (freckles):

Do you see 3 requirements?

OpenStudy (freckles):

And remember you only care about satisfying the if part to see Rolle's theorem applies.

OpenStudy (freckles):

that first line "If f" <-- not "If I"

OpenStudy (anonymous):

So f(c) must equal 0 at some point on the function?

OpenStudy (freckles):

Only look at the if part to see what is required to apply the theorem

OpenStudy (freckles):

requirement number 1) f is continuous on [a,b]

OpenStudy (freckles):

requirement number 2) is?

OpenStudy (anonymous):

Differentiable on open interval.

OpenStudy (anonymous):

And so the 3rd requirement is the function must = 0 at some point? Correct me if I'm wrong...

OpenStudy (freckles):

f(a)=f(b)

OpenStudy (freckles):

f(a)=f(b) is all in the if part

OpenStudy (freckles):

also not all

OpenStudy (freckles):

For example say we have f(x)=x^2. f is continuous on [-3,3] f is differntiable on (-3,3) f(3)=f(-3) So we can apply rolle's theorem That is there is at least one value c in the interval (-3,-3) such that f'(c)=0 Well f'(x)=2x So 2x=0 when x=0 0 is indeed in between -3 and 3

OpenStudy (anonymous):

oh okay... :) Thanks!

OpenStudy (anonymous):

do you need help with anything?

OpenStudy (freckles):

Nah I'm cool. Thanks for asking though :)